1. -4x-5y=-2
y=-14-8x

Answer by mananth(16949)

-4 x -5 y = -2 .............1
8 x + 1 y = -14 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 5
-4x-5y = -2
40x+5y = -70
Add the two equations
36 x = -72.00
/ 36
x = -2
plug value of x in (1)
-4 x + -5 y = -2
8 + -5 y = -2
-5 y = -2
-5 y = -10
y = 2